### v1

@@ 4.4 Example of Non-monotone DP
    $G_{t}$ is edgeless, so the only inclusion-wise minimal vertex cover is an empty set. Let $C_{\emptyset}=0^{\omega}$ be a tuple that corresponds to $\emptyset$. Then
@@
    Observe that a set $S$ is an inclusion-wise minimal vertex cover in $G_{t}$ if and only if it is a union of inclusion-wise minimal vertex covers $S_{1}$ in $G_{t_{1}}$ and $S_{2}$ in $G_{t_{2}}$.
    Recall that for $C_{1}=(x_{1},\ldots,x_{\omega}),\ C_{2}=(y_{1},\ldots,y_{\omega})$ we denote by $C_{1}+C_{2}$ the tuple ${(x_{1}+y_{1},\ldots,x_{\omega}+y_{\omega})}$. Then
    $\mathcal{F}_{t}=\{C_{1}+C_{2}|C_{1}\in\mathcal{F}_{t_{1}},C_{2}\in\mathcal{F}_{t_{2}}\}.$
@@
    $\mathcal{F}_{t}=\{\rho_{a\rightarrow b}(C^{\prime})|C^{\prime}\in\mathcal{F}_{t}^{\prime}\},$
@@
    Observe that a set $S$ is an inclusion-wise minimal vertex cover in $G_{t}$ if and only if it is a union of an inclusion-wise minimal vertex cover $S^{\prime}$ in $G_{t^{\prime}}$ and an inclusion-wise minimal set of vertices that covers new edges, which is either the set of all vertices of color $a$ or the set of all vertices of color $b$.
    Recall that $n_{i}$ denotes the number of vertices of color $i$ in $G_{t}$, and we can compute it by induction for each $i\in[\omega]$ for each $t\in V(\mathcal{D})$ in time $|V(\mathcal{D})|$.
@@
    $\mathcal{F}_{t}=\{all_{a}(C^{\prime})|C^{\prime}\in\mathcal{F}_{t^{\prime}}\}\cup\{all_{b}(C^{\prime})|C^{\prime}\in\mathcal{F}_{t^{\prime}}\}.$
The core corresponding to this dynamic programming is clearly not monotone. The value of any $(\mathfrak{C},\mathcal{D})$-witness $\alpha$ on the leaves of $\mathcal{D}$ is equal to $\{C_{\emptyset}\}$, so for any membership function $\rho$ either $S_{\rho}(\mathcal{D},\alpha)=\emptyset$ or $S_{\rho}(\mathcal{D},\alpha)=V(G)$.
@@
Observe that we can also solve $k$-Min Vertex Cover by running the same dynamic programming as for $k$-Vertex Cover and then choosing the solution of the smallest size among the found ones. The worst case running time is asymptotically the same, so we do not benefit from dropping the monotonicity constraint.
@@ 5 MSO1 Has Monotone DP
Given a cliquewidth decomposition $\mathcal{D}$ of a graph $G$, an $MSO_{1}$ formula $\Phi$ can be evaluated on $G$ in time $|V(\mathcal{D})|\cdot c$, where $c$ is a constant depending on $\Phi$ and the width of $\mathcal{D}$.
The proof of Theorem 5.1 will give us a dynamic programming over the decomposition of $G$. We will show that for vertex problems this dynamic programming is monotone, which, combined with Theorem 4.6, will give us an FPT algorithm for diverse vertex problems expressible in $MSO_{1}$.
5.1 Proof of the Courcelle’s Theorem for Cliquewidth Using Reduced Evaluation Trees
Let $\Phi$ be an $MSO_{1}$ formula in the prenex normal form and let $\Psi(z_{1},\ldots,z_{q})$ be the quantifier-free part of $\Phi$. Let $q_{v}$ be the number of variables quantified over single vertices (called individual variables) and $q_{s}$ be the number of variables quantified over sets of vertices (called set variables). The total quantifier depth of $\Phi$ is $q=q_{v}+q_{s}$
@@ 5.1 Proof of the Courcelle’s Theorem for Cliquewidth Using Reduced Evaluation Trees
Evaluation tree for a formula $\Phi$ and graph $G$ is a rooted tree $F$ with $q+1$ levels constructed as follows. For each level $\ell\in[q]$ a node $w$ of the $\ell$-th level corresponds to a particular choice of the first $\ell$ variables. If $z_{\ell+1}$ is an individual variable, then $w$ has $n$ children, representing all possible choices of vertices for $z_{\ell+1}$. If $z_{\ell+1}$ is a set variable, then $w$ has $2^{n}$ children, representing all possible choices of subsets of vertices for $z_{\ell+1}$.
The answer whether $\Phi$ is satisfied on $G$ can be obtained by traversing $F$ bottom-up. The value of each internal node depends on the values of its children. If a node $w$ corresponds to a universally quantified variable, it must have all children satisfied to be satisfied, if it corresponds to an existentially quantified variable, it must have at least one satisfied child to be satisfied. However, $F$ can be large. But it turns out that its size can be reduced.
@@
When a set assigned to a variable $z_{i}$ contains a vertex $u$ that is not assigned to any individual variable, then u does not participate in any of the membership predicates. Hence, the evaluation of $\Psi$ is independent on its presence in $z_{i}$.
The reduction goes inductively bottom-up. Each leaf of $F$ corresponds to an evaluation of $\Psi$ on a subgraph $G^{\prime}$ of $G$ induced by $q_{v}$ vertices assigned to individual variables. The number of non-isomorphic graphs on $q_{v}$ vertices is bounded by $2^{q_{v}\choose 2}$. The number of possible assignments of vertices of such graph to individual variables is bounded by ${q_{v}}^{q_{v}}$ and the number of possible assignments of sets of vertices to set variables is bounded by $(2^{q_{v}})^{q_{s}}$. In total, we get at most $2^{\mathcal{O}(q_{v}q)}$ non-isomorphic combinations of $G^{\prime}$ and variable assignments. These combinations are isomorphism classes of the leaves. Since leaves from the same isomorphism class have the same value in the evaluation, the evaluation of node $w$ of level $q$ is independent of whether there is one or more leaf from the same class among its children, so it suffices to keep only one leaf for each isomorphism class that appear there. After this reduction each node of level $q$ has at most $2^{\mathcal{O}(q_{v}q)}$ children.
Now we can distinguish nodes of level $q-1$ by sets of their children. There are at most $2^{2^{\mathcal{O}(q_{v}q)}}$ possible sets of children. Each of these sets defines the isomorphism class of trees of height one. Now reduce isomorphic subtrees under the same parent analogously as leaves.
This reduction procedure can be repeated consecutively for levels $q-2,\ldots,1$.
@@
Now we will show that given a cliquewidth decomposition $\mathcal{D}$ of $G$ of width $k$ we can construct $F^{-}$ without constructing $F$.
@@
Partial evaluation tree for a formula $\Phi$ and a node $t$ of the decomposition $\mathcal{D}$ of $G$ is a tree obtained from the evaluation tree for $\Phi$ and $G_{t}$ by adding to each node $w$ of level $\ell$ where $z_{\ell+1}$ is an individual variable an extra child labeled by $ext$ representing the case where $z_{\ell+1}$ does not belong to $G_{t}$. We say the value of $z_{\ell}$ is external.
@@
Let $\Phi(z_{1},\ldots,z_{\ell})$ be an $MSO_{1}$ formula in a prenex normal form, where $z_{1},\ldots,z_{\ell}$ are free variables and $z_{\ell+1},\ldots,z_{q}$ are quantified variables. Let $F_{1},F_{2}$ be partial evaluation trees for $\Phi(z_{1},\ldots,z_{\ell})$ and $G_{t_{1}}$ and $G_{t_{2}}$ respectively, where $G_{t_{1}}$ and $G_{t_{2}}$ are disjoint, $u$ is the root of $F_{1}$ and $w$ is the root of $F_{2}$. Product of $F_{1},F_{2}$ is a partial evaluation tree $F\coloneq F_{1}\otimes F_{2}$ defined recursively.
  • If the height of $F_{1},F_{2}$ is 00 (that means $\ell=q$), then $F$ is a tree of height 00 with root $(u,w)$.
  • If the height of $F_{1},F_{2}$ greater than 00
@@
    • If $z_{\ell+1}$ is an individual variable, $F_{1}$ has children subtrees $F_{1}^{0},F_{1}^{1},\ldots,F_{1}^{n}$ labeled by $ext,s_{1},\ldots,s_{n}\in V(G_{t_{1}})$, $F_{2}$ has children subtrees $F_{2}^{0},F_{2}^{1},\ldots,F_{2}^{m}$ labeled by $ext,p_{1},\ldots,p_{m}\in V(G_{t_{2}})$, then $F$ is a tree with root $(u,v)$ which has: for each $i\in[n]$ a child subtree $F_{1}^{i}\otimes F_{2}^{0}$ labeled by $s_{i}$, for each $j\in[m]$ a child subtree $F_{0}^{i}\otimes F_{2}^{j}$ labeled by $p_{j}$ and a child subtree $F_{1}^{0}\otimes F_{2}^{0}$ labeled by $ext$.
@@
We will show that we can reduce partial evaluation trees bottom up, similarly to regular evaluation trees. Being able to perform this reduction after each step of the construction, we will keep the size of partial evaluation trees for the inner nodes of the decomposition small and make the construction efficient.
We will start with defining the isomorphism classes of the leaves.
@@
Let $\mathcal{D}$ be a cliquewidth decomposition of width $k$, $F_{t}$ be a partial evaluation tree for a formula $\Phi$ and $t\in V(\mathcal{D})$. A configuration of a node $u$ of $F_{t}$ of level $\ell$ is a combination of:
@@
  • a coloring of these vertices $\mu:V(G^{\prime})\rightarrow[k]$,
  • an assignment of values from $V(G^{\prime})\cup\{ext\}$ to the individual variables among the first $\ell$ variables,
  • an assignment of subsets of $V(G^{\prime})$ to set variables among the first $\ell$ variables.
For the leaves of $F_{t}$ (nodes of level $q$) there are at most $2^{q_{v}\choose 2}\cdot k^{q_{v}}\cdot(q_{v}+1)^{q_{v}}\cdot(2^{q_{v}})^{q_{s}}=2^{\mathcal{O}(q_{v}(q+\log k))}$ different configurations. These configurations define isomorphism classes of the leaves (subtrees of height 00 in $F$). For each $h\in[q]$, the isomorphism classes of the subtrees of height $h$ are defined by the sets of isomorphism classes of their children.
Performing the same reduction as for regular evaluation tree we get reduced partial evaluation tree $F_{t}^{-}$.
@@
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q+1$.}$
@@
If we prune in $F_{t}^{-}$ all configurations containing external choices, we get an evaluation tree isomorphic to the full evaluation tree for $\Phi$ and $G_{t}$.
In order to make the reduction procedure deterministic, for each isomorphism class of trees of height $h$ we can fix one representant. Then $reduce(F)$ replaces $F$ by the representant of its isomorphism class.
It remains to prove that we can perform this reduction after each step of the construction of a partial evaluation tree for $G$.
@@
Let $t_{1},t_{2}\in V(\mathcal{D})$ such that $G_{t_{1}},G_{t_{2}}$ are disjoint, $F_{1},F_{1}^{\prime}$ be partial evaluation trees for $\Phi$ and $t_{1}$, $F_{2},F_{2}^{\prime}$ be partial evaluation trees for $\Phi$ and $t_{2}$, $T_{1}\sim T_{1}^{\prime}$ be subtrees of $F_{1},F_{1}^{\prime}$, respectively, rooted at level $\ell$, $T_{2}\sim T_{2}^{\prime}$ be subtrees of $F_{2},F_{2}^{\prime}$, respectively, rooted at level $\ell$. Then ${T_{1}\otimes T_{2}\sim T_{1}^{\prime}\otimes T_{2}^{\prime}}$.
In order to prove Lemma 5.12, we will firstly introduce the notion of equivalent paths and prove similar fact for combinations of such paths.
Definition 5.13 (Equivalent paths).
Let $F,F^{\prime}$ be partial evaluation trees for $\Phi$ and $t\in V(\mathcal{D})$. Two paths from the root $P\in F$, $P^{\prime}\in F^{\prime}$ of the same length are equivalent if the configurations of their endpoints are the same.
Observation 5.14.
If two paths $P,P^{\prime}$ of length $\ell$ are equivalent, then for each $i\in[\ell]$ their prefixes of length $i$ are also equivalent.
Observation 5.15.
If two paths $P,P^{\prime}$ of length $\ell$ are equivalent, then for each $i\in[\ell]$ the $i$-th node of $P$ is labeled by $ext$ if and only if the $i$-th node of $P^{\prime}$ is labeled by $ext$.
Definition 5.16 (Compatible paths).
Let $F_{1},F_{2}$ be partial evaluation trees for $\Phi$ and $t_{1},t_{2}\in V(\mathcal{D})$ respectively, where $G_{t_{1}},G_{t_{2}}$ are disjoint. Two paths from the root $P_{1}\in F_{1}$, $P_{2}\in F_{2}$ of the same length $\ell$ are compatible if for each $i\in[\ell]$ such that $z_{i}$ is an individual variable at most one of the labels of the $i$-th node of $P_{1}$ and $P_{2}$ is not $ext$.
Definition 5.17 (Combination of paths).
For two compatible paths $P_{1},P_{2}$ of length $\ell$ their combination is the path $P\coloneq P_{1}\otimes P_{2}$ of the same length such that for each $i\in[\ell]$ it holds:
  • if $z_{i}$ is a set variable, the label of the $i$-th node of $P_{1}$ is $S_{1}$ and the label of the $i$-th node of $P_{2}$ is $S_{2}$, then the label of the $i$-th node of $P$ is $S_{1}\cup S_{2}$,
  • if $z_{i}$ is an individual variable and the label of the $i$-th node of $P_{1}$ is not $ext$, then the label of the $i$-th node of $P$ is equal to the label of the $i$-th node of $P_{1}$, otherwise it is equal to the label of the $i$-th node of $P_{2}$.
Observation 5.18.
When we construct the tree product $F\coloneq F_{1}\otimes F_{2}$, we combine these subtrees $T_{1}\subseteq F_{1},\ T_{2}\subseteq F_{2}$ whose paths to the root of $F_{1}$ and $F_{2}$ respectively are compatible. Moreover, the path of $T_{1}\otimes T_{1}$ to the root of $F$ is the combination of these two paths.
Lemma 5.19 (Combining equivalent paths).
Let $t_{1},t_{2}\in V(\mathcal{D})$ such that $G_{t_{1}},G_{t_{2}}$ are disjoint, $F_{1},F_{1}^{\prime}$ be partial evaluation trees for $\Phi$ and $t_{1}$, $F_{2},F_{2}^{\prime}$ be partial evaluation trees for $\Phi$ and $t_{2}$. If path $P_{1}\in F_{1}$ is equivalent to path $P_{1}^{\prime}\in F_{1}^{\prime}$ and path $P_{2}\in F_{2}$ is equivalent to path $P_{2}^{\prime}\in F_{2}^{\prime}$, $P_{1}$ and $P_{2}$ are compatible, then $P_{1}^{\prime}$ and $P_{2}^{\prime}$ are also compatible and $P_{1}\otimes P_{2}$ is equivalent to $P_{1}^{\prime}\otimes P_{2}^{\prime}$.

### v2

@@ 4.4 Example of Non-monotone DP
    $G_{t}$ is edgeless, so the only inclusion-wise minimal vertex cover is an empty set. Let $C_{\emptyset}$ be a tuple that corresponds to a vertex cover $\emptyset$. Then
@@
    Observe that a set $S$ is an inclusion-wise minimal vertex cover of $G_{t}$ if and only if it is a union of inclusion-wise minimal vertex covers $S_{1}$ of $G_{t_{1}}$ and $S_{2}$ of $G_{t_{2}}$. Recall that for $C_{1}=(x_{1},\ldots,x_{\omega}),\ C_{2}=(y_{1},\ldots,y_{\omega})$ we denote by $C_{1}+C_{2}$ the tuple $(x_{1}+y_{1},\ldots,x_{\omega}+y_{\omega})$}. Then
    $\mathcal{F}_{t}=\{C_{1}+C_{2}:C_{1}\in\mathcal{F}_{t_{1}},C_{2}\in\mathcal{F}_{t_{2}},|C_{1}+C_{2}|\leq k\}.$
@@
    $\mathcal{F}_{t}=\{\rho_{a\rightarrow b}(C^{\prime}):C^{\prime}\in\mathcal{F}_{t}^{\prime}\},$
@@
    Observe that a set $S$ is an inclusion-wise minimal vertex cover in $G_{t}$ if and only if it is a union of an inclusion-wise minimal vertex cover $S^{\prime}$ in $G_{t^{\prime}}$ and an inclusion-wise minimal set of vertices that covers new edges, which is either the set of all vertices of color $a$ or the set of all vertices of color $b$. Recall that $n_{i}$ denotes the number of vertices of color $i$ in $G_{t}$, and we can compute it by induction for each $i\in[\omega]$ for each $t\in V(\mathcal{D})$ in time $|V(\mathcal{D})|$.
@@
    $\mathcal{F}_{t}=\{all_{a}(C^{\prime}):C^{\prime}\in\mathcal{F}_{t^{\prime}}\}\cup\{all_{b}(C^{\prime}):C^{\prime}\in\mathcal{F}_{t^{\prime}}\}.$
The core corresponding to this dynamic programming is clearly not monotone. The value of any $(\mathfrak{C},\mathcal{D})$-witness $\alpha$ on the leaves of $\mathcal{D}$ is equal to $C_{\emptyset}$, so for any membership function $\rho$ either $S_{\rho}(\mathcal{D},\alpha)=\emptyset$ or $S_{\rho}(\mathcal{D},\alpha)=V(G)$.
@@
Observe that we can also solve $k$-Min Vertex Cover by running the same dynamic programming as for $k$-Vertex Cover and then choosing the solution of the smallest size among the found ones. The worst-case running time is asymptotically the same, so we do not benefit from dropping the monotonicity constraint.
@@ 5 MSO1 Has Monotone DP
Given a cliquewidth decomposition $\mathcal{D}$ of a graph $G$, an $MSO_{1}$ formula $\Phi$ can be evaluated on $G$ in time $c\cdot|V(\mathcal{D})|$, where $c$ is a constant depending only on $\Phi$ and the width of $\mathcal{D}$.
The proof of Theorem˜5.1 will give us dynamic programming over the decomposition of $G$. In Section˜5.2, we will show that for vertex problems, this dynamic programming is monotone, which, combined with Theorem˜3.6, gives us Corollary˜1.5.
5.1 Proof of Courcelle’s Theorem for Cliquewidth Using Reduced Evaluation Trees
We prove Theorem˜5.1 using partial evaluation trees. Let $\Phi$ be an $MSO_{1}$ formula in the prenex normal form and let $\Psi(z_{1},\ldots,z_{q})$ be the quantifier-free part of $\Phi$. Let $q_{v}$ be the number of variables quantified over single vertices (called individual variables) and $q_{s}$ be the number of variables quantified over sets of vertices (called set variables). The total quantifier depth of $\Phi$ is $q=q_{v}+q_{s}$
@@ 5.1 Proof of Courcelle’s Theorem for Cliquewidth Using Reduced Evaluation Trees
Evaluation tree for a formula $\Phi$ and graph $G$ is a rooted tree $F$ with $q+1$ levels (where 0 is the level of the root) constructed as follows. For each level $\ell\in[q]$, each node $w$ of the $\ell$-th level corresponds to a particular choice of the first $\ell$ variables. If $z_{\ell+1}$ is an individual variable, then $w$ has $n$ children, representing all possible choices of vertices for $z_{\ell+1}$. If $z_{\ell+1}$ is a set variable, then $w$ has $2^{n}$ children, representing all possible choices of subsets of vertices for $z_{\ell+1}$.
The answer to whether $\Phi$ is satisfied on $G$ can be obtained by traversing $F$ bottom-up. The value of a leaf is true if the valuation corresponding to this leaf satisfies $\Phi$. The value of each internal node depends on the values of its children. If a node $w$ corresponds to a universally quantified variable, it must have all children satisfied to be satisfied, if it corresponds to an existentially quantified variable, it must have at least one satisfied child to be satisfied. However, $F$ can be large. But it turns out that its size can be reduced.
@@
When a set assigned to a variable $z_{i}$ contains a vertex $u$ that is not assigned to any individual variable, then $u$ does not participate in any of the membership predicates. Hence, the evaluation of $\Psi$ is independent of its presence in $z_{i}$.
The reduction goes inductively bottom-up. Each leaf of $F$ corresponds to an evaluation of $\Psi$ on a subgraph $G^{\prime}$ of $G$ induced by $q_{v}$ vertices assigned to individual variables. The number of non-isomorphic graphs on $q_{v}$ vertices is bounded by $2^{\binom{q_{v}}{2}}$. The number of possible assignments of vertices of such a graph to individual variables is bounded by ${q_{v}}^{q_{v}}$, and the number of possible assignments of sets of vertices to set variables is bounded by $(2^{q_{v}})^{q_{s}}$. In total, we get at most $2^{\mathcal{O}(q_{v}q)}$ non-isomorphic combinations of $G^{\prime}$ and variable assignments. These combinations define isomorphism classes of the leaves. Since leaves from the same isomorphism class have the same value in the evaluation, the evaluation of node $w$ of level $q$ is independent of whether there are one or more leaves from the same class among its children, so it suffices to keep only one leaf for each isomorphism class that appears there. After this reduction, each node of level $q$ has at most $2^{\mathcal{O}(q_{v}q)}$ children. Now we can distinguish nodes of level q1q-1 by sets of their children. There are at most $2^{2^{\mathcal{O}(q_{v}q)}}$ possible sets of children. Each of these sets defines the isomorphism class of trees of height one. Now reduce isomorphic subtrees under the same parent analogously as leaves. This reduction procedure can be repeated consecutively for levels $q-2,\ldots,1$.
@@
Now we will show that given a cliquewidth decomposition $\mathcal{D}$ of $G$ of width $\omega$, we can construct $F^{-}$ without constructing $F$.
@@
Partial evaluation tree for a formula $\Phi$ and a node $t$ of the decomposition $\mathcal{D}$ of $G$ is an evaluation tree for $\Phi$ and $G_{t}$ where for each individual variable we additionally consider a choice of external value (not belonging to $G_{t}$) for this variable, and add an extra branch labeled by extext representing this choice. That means for each $\ell$ such that $z_{\ell+1}$ is an individual variable each node of level $\ell$ has $|V(G_{t})|+1$ children.
@@
Let $\Phi(z_{1},\ldots,z_{\ell})$ be an $MSO_{1}$ formula in a prenex normal form, where $z_{1},\ldots,z_{\ell}$ are free variables and $z_{\ell+1},\ldots,z_{q}$ are quantified variables. Let $F_{1},F_{2}$ be partial evaluation trees for $\Phi(z_{1},\ldots,z_{\ell})$ and $G_{t_{1}}$ and $G_{t_{2}}$ respectively, $u$ is the root of $F_{1}$ and $w$ is the root of $F_{2}$. Product of $F_{1},F_{2}$ is a partial evaluation tree $F\coloneq F_{1}\otimes F_{2}$ defined recursively.
  • If the height of $F_{1},F_{2}$ is 0 (that means $\ell=q$), then $F$ is a tree of height 0 with root $(u,w)$.
  • If the height of $F_{1},F_{2}$ greater than 0
@@
    • If $z_{\ell+1}$ is an individual variable, $F_{1}$ has children subtrees $F_{1}^{0},F_{1}^{1},\ldots,F_{1}^{n}$ labeled by $ext,s_{1},\ldots,s_{n}\in V(G_{t_{1}})$, $F_{2}$ has children subtrees $F_{2}^{0},F_{2}^{1},\ldots,F_{2}^{m}$ labeled by $ext,p_{1},\ldots,p_{m}\in V(G_{t_{2}})$, then $F$ is a tree with root $(u,v)$ which has: for each $i\in[n]$ a child subtree $F_{1}^{i}\otimes F_{2}^{0}$ labeled by $s_{i}$, for each $j\in[m]$ a child subtree $F_{1}^{0}\otimes F_{2}^{j}$ labeled by $p_{j}$ and a child subtree $F_{1}^{0}\otimes F_{2}^{0}$ labeled by extext.
@@
We will show that we can reduce partial evaluation trees bottom up, similarly to regular evaluation trees. By being able to perform this reduction after each step of the construction, we will keep the size of partial evaluation trees for the inner nodes of the decomposition small and make the construction efficient.
We start by defining the isomorphism classes of the leaves.
@@
Let $\mathcal{D}$ be a cliquewidth decomposition of width $\omega$, $F_{t}$ be a partial evaluation tree for a formula $\Phi$, and $t\in V(\mathcal{D})$. A configuration of a node $u$ of $F_{t}$ of level $\ell$ is a combination of:
@@
  • a coloring of these vertices $\mu:V(G_{t}^{\prime})\rightarrow[\omega]$,
  • an assignment $\rho_{I}$ of values from $V(G_{t}^{\prime})\cup\{ext\}$ to the individual variables among the first $\ell$ variables,
  • an assignment $\rho_{S}$ of subsets of $V(G_{t}^{\prime})$ to set variables among the first $\ell$ variables.
For the leaves of $F_{t}$ (nodes of level $q$) there are at most $2^{\binom{q_{v}}{2}}\cdot\omega^{q_{v}}\cdot(q_{v}+1)^{q_{v}}\cdot(2^{q_{v}})^{q_{s}}=2^{\mathcal{O}(q_{v}(q+\log\omega))}$ different configurations. These configurations define isomorphism classes of the leaves (subtrees of height 0 in $F_{t}$). For each $h\in[q]$, the isomorphism classes of the subtrees of height $h$ are defined by the sets of isomorphism classes of their children.
Performing the same reduction as for the regular evaluation tree, we get reduced partial evaluation tree $F_{t}^{-}$.
@@
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log\omega))}}}}}\end{array}\right\}\text{\scriptsize$q+1$.}$
@@
If we prune in $F_{t}^{-}$ all branches containing external choices, we get an evaluation tree isomorphic to the full evaluation tree for $\Phi$ and $G_{t}$.
In order to make the reduction procedure deterministic, for each isomorphism class of trees of height $h$, we can fix one representative. Then $reduce(F)$ replaces $F$ by the representative of its isomorphism class.
It remains to be proven that we can perform this reduction after each step of the construction of a partial evaluation tree for $G$.
Lemma 5.12.
We start by a few observations.
Observation 5.13.
Let $F_{1},F_{2}$ be partial evaluation trees for $\Phi$, $u,v$ be nodes of level $\ell>0$ of $F_{1}$ and $F_{2}$, respectively. If the configurations of $u$ and $v$ are the same, then the configurations of their parents are the same.
Observation 5.14.
Let $F_{1},F_{2}$ be partial evaluation trees for $\Phi$, $T_{1},T_{2}$ be subtrees of $F_{1}$ and $F_{2}$ rooted at the same level. Then if $T_{1}\sim T_{2}$, the configurations of their roots are the same.
We say two configurations of nodes of level $\ell$ are compatible if for each $i\leq\ell$, where $z_{i}$ is individual variable, the value assigned to $z_{i}$ is external in at least one of these configurations. Observe that while computing the tree product $F_{1}\otimes F_{2}$ recursively, we combine exactly these subtrees of $F_{1},F_{2}$ rooted at the same level whose root configurations are compatible.
Definition 5.15 (Union of configurations).
Let $(G_{1},\mu_{1},\rho_{I}^{1},\rho_{S}^{1}),(G_{2},\mu_{2},\rho_{I}^{2},\rho_{S}^{2})$ be a pair of compatible configurations of nodes of level $\ell$ of partial evaluation trees for the same formula $\Phi$. Their union is a configuration $(G,\mu,\rho_{I},\rho_{S})$, where:
  • $G=G_{1}\uplus G_{2}$, i.e. $G$ is a disjoint union of $G_{1}$ and $G_{2}$,
  • $\mu(v)=\begin{cases}\mu_{1}(v)&\text{ if }v\in V(G_{1}),\\ \mu_{2}(v)&\text{ otherwise,}\end{cases}$
  • $\rho_{I}=\begin{cases}\rho_{I}^{1}(z_{i})&\text{ if }\rho_{I}^{1}(z_{i})\neq ext,\\ \rho_{I}^{2}(z_{i})&\text{ otherwise,}\end{cases}$
  • $\rho_{S}(z_{i})=\rho_{S}^{1}(z_{i})\uplus\rho_{S}^{2}(z_{i})$.
Observation 5.16.
Let $F_{1},F_{2}$ be partial evaluation trees for $\Phi$ and $t_{1},t_{2}$, respectively and $T_{1},T_{2}$ be subtrees of $F_{1},F_{2}$ rooted at the same level. Then $T_{1}\otimes T_{2}$ is a subtree of $F_{1}\otimes F_{2}$ if and only if the root configurations of $T_{1}$ and $T_{2}$ are compatible. Moreover, if $G_{t_{1}}$ and $G_{t_{2}}$ are disjoint, the root configuration of $T_{1}\otimes T_{2}$ in $F_{1}\otimes F_{2}$ is a union of the root configurations of $T_{1}$ and $T_{2}$.
Lemma˜5.12 is a direct application of the following lemma, with $\ell=0$.
@@
Let $t_{1},t_{2}\in V(\mathcal{D})$ such that $G_{t_{1}},G_{t_{2}}$ are disjoint, $F_{1},F_{1}^{\prime}$ be partial evaluation trees for $\Phi$ and $t_{1}$, $F_{2},F_{2}^{\prime}$ be partial evaluation trees for $\Phi$ and $t_{2}$, $T_{1}\sim T_{1}^{\prime}$ be subtrees of $F_{1},F_{1}^{\prime}$, respectively, rooted at level $\ell$, $T_{2}\sim T_{2}^{\prime}$ be subtrees of $F_{2},F_{2}^{\prime}$, respectively, rooted at level $\ell$. Then, if the root configurations of $T_{1}$ and $T_{2}$ are compatible, $T_{1}\otimes T_{2}$ is a subtree of $F_{1}\otimes F_{2}$, $T_{1}^{\prime}\otimes T_{2}^{\prime}$ is a subtree of $F_{1}^{\prime}\otimes F_{2}^{\prime}$ and moreover $T_{1}\otimes T_{2}\sim T_{1}^{\prime}\otimes T_{2}^{\prime}$}.